T 10:15 P.M., a train left Petersburg, Virginia, travelling east at 40 mi/h. Two hours later, another train left the same station travelling west at 35 mi/h. At what time will the trains be 230 mi apart?
Answer by mananth(16949)

towards east train - 10:15 ----40 mph
Towards Weat -----35 mph----12:15
in 2 hours East bound train would have traveled 80 miles before west bound train starts.
Distance they to be apart = 230.
230-80 = 150 miles is the distance to be covered.
add up speeds 35+40 =75 mph
t=d/r
150/75 = 2 hours
they will be 230 miles apart at 2:15 pm.( 2hours after west bound trains leaves)